# Computable Følner sequences of amenable groups

### v1

@@ Abstract
The paper considers computable Følner sequences in computably enumerable amenable groups, We extend some basic results of M. Cavaleri on existence of such sequences to the case of groups where finite generation is not assumed. We also initiate some new directions in this topic, for example complexity of families of effective Følner sequences. Possible extensions of this approach to metric groups are also discussed.
@@ 1 Introduction
Since being finitely generated is not necessary for amenability, the question arises what happens if we consider the case of recursively presented groups without the assumption of finite generation. According the approach of computable algebra, the question concerns the class of computably enumerable numbered groups and the subclass of computable numbered groups, a counterpart of decidability of the word problem. These notions are thoroughly discribed in Section 2. The following theorem generalizes aforementioned results of Cavaleri to the case of computably enumerable numbered groups.
@@
We also compare convergence moduli of sequences of means corresponding to these Følner sequences. In particular we show that they cannot be estimated by primitive recursive functions.
@@
The author is grateful to M. Cavaleri and T. Ceccherini-Silberstein for reading of some preliminary versions of the paper and helpful remarks.
@@ 2 Preliminaries
Concerning other details, let us mention that we often identify finite sets $F\subset\mathbb{N}$ with their Gödel numbers. For any sets $X$ and $Y$ we will write $X\subset_{fin}Y$ to denote that $X$ is a finite subset of $Y$. For any $i\in\mathbb{N}$, we denote the set $\{1,2,\ldots,i\}$ by $[i]$. Throughout this paper a group (metric group) $G$ is a countable/separable group without any presumption about its generating set.
@@ 2.1 Computabie presentations
We emphasize that this group is computable! For every recursively presented group $G=\langle X\rangle$ and a natural homomorphism $\rho:\mathbb{F}_{\omega}\rightarrow G$ (taking $\omega$ onto $X$) we obtain that $\rho$ is a numbering of a computably enumerable presentation of $G$. This argument can be generalized as follows.
@@
    The set
    $\mathsf{MultT}:=\{(i,j,k):\quad\nu(i)\nu(j)=\nu(k)\}$
    is computable if and only if the set $\{(n_{1},n_{2}):\nu(n_{1})=\nu(n_{2})\}$ is computable;
@@
From now on we consider computably enumerable groups together with presentations satisfying the conditions of Lemma 2.3.
@@ 2.2 Amenability
$\forall x\in D\hskip 14.22636pt\frac{|F\setminus xF|}{|F|}\leq\frac{1}{n}$ (1)
@@
A sequence $(F_{j})_{j\in\mathbb{N}}$ of non-empty finite subsets of $G$ is a Følner sequence if for every $g\in G$ the following condition holds:
@@
Suppose that $G$ satisfies the Følner conditions. Let $(F_{j}\,|\,j\in J)$ be a left Følner’s net of $G$. Consider, for each $j\in J$, the mean with finite support $m_{j}:\ell^{\infty}(G)\to\mathbb{R}$ defined by
@@
The following facts are well known and/or easy to prove.
@@ 3 Effective amenability of computably enumerable groups
In this section $(G,\nu)$ is a numbered metric group such that $\nu$ is a homomorphism from $(\mathbb{N},*,^{-1},1)$ to $G$. In the case of the Følner condition of amenability, we consider two types of effectiveness.
@@
Some variants of these notions were characterized by M. Cavaleri in [4] in the case of finitely generated groups. The goal of this section is an adaptation of these characterizations in our general case.
@@ 3.1 Computable Reiter’s functions
The main result of this part, Theorem 3.5, is a natural generalization of a theorem of M. Cavaleri from [4] (Theorem 3.1) to the case of groups which are not finitely generated. Although we use the same arguments, its adaption to our case is not straightforward. Throughout this section we assume that $(G,\nu)$ is a computably enumerable group under a homomrphism $\nu$ as in the beginning of the section.
The following notation will be used below. If $f:\mathbb{N}\to\mathbb{R}^{+}$ is summable and $g\in G$, then let
@@
Having $x\in\mathbb{N}$ and an equivalence relation $E$ on the set $F\cup x^{-1}*F$ fix a set $F_{0}\subset F$ of representatives of the classes of $E$ in $F$. Then we define the positive rational number:
$M_{E,F}^{x}(f):=\frac{\sum\{|(E_{F,1}(f)(v)-E_{F,1}(f)(x^{-1}*v))|\,|\,v\in F_{0}\}}{\sum_{v\in F}f(v)}.$
We denote by $P$ the canonical equivalence relation of the set $F\cup x^{-1}*F$, i.e. the partition into sets $\{\nu^{-1}(\nu(k))\,|\,k\in F\cup x^{-1}*F\}$. Then for every $x\in\mathbb{N}$ we have
@@
For any two equivalence relations $E$ and $E^{\prime}$ of the set $F\cup x^{-1}*F$ with $E\subseteq E^{\prime}$ the triangle inequality implies $M_{E,F}^{x}(f)\geq M_{E^{\prime},F}^{x}(f)$. In particular, for any $x\in\mathbb{N}$ and $E\subseteq P$ on $F\cup x^{-1}*F$ we have:
@@
The algorithm $\varkappa(n,D,f)$ starts as follows. For an input $f$ let $F=\mathsf{supp}(f)$. Put $P_{0}$ to be the (finest) partition of $\bigcup_{g\in D}g^{-1}*F$ into singletons. At the $m$-th step of the enumeration of $\{(n_{1},n_{2}):\nu(n_{1})=\nu(n_{2})\}$ we are trying to merge classes of the equivalence relation $P_{m-1}$ on $\bigcup_{g\in D}g^{-1}*F$ already constructed at step m1m-1. We do so when we meet $(n_{1},n_{2})\not\in P_{m-1}$ with $\nu(n_{1})=\nu(n_{2})$. In this case we just merge the classes of $n_{1}$ and $n_{2}$. In this way we obtain $P_{m}\subseteq P$. Then we verify if $M_{P_{m},F}^{x}(f)\leq\frac{1}{n}$ for all $x\in D$. We stop $\varkappa(n,D,f)$ when these inequalities hold. In this case by (5) and (6) the function $\nu_{G,1}(f)$ is $\frac{1}{n}$-invariant. If there exist $x$, such that $M_{P_{m},F}^{x}(f)>\frac{1}{n}$ and $P_{m}=P$, then the function $\nu_{G,1}(f)$ is not $\frac{1}{n}$-invariant. Note that there is no algorithm for recognizing the latter possibility.
@@ 3.2 Effective amenability of computable groups
Fix $n_{1},n_{2}$. Let $E$ be the set $\{n_{1},n_{2}\}$. We use the algorithm for (ii) to find a set $F$ corresponding to $3$ and $E$, i.e. $\nu(F)\in F\o l_{G,\nu(E)}(3)$ and $|F|=|\nu(F)|$. Let $F=\{f_{1},f_{2},\ldots,f_{k}\}$.
For each $i\in\{1,2\}$ we define $\Sigma_{i}\subseteq\{(f,f^{\prime})\,|\,\nu(n_{i})\nu(f)=\nu(f^{\prime})\,,\,f,f^{\prime}\in F\}$ by the following procedure. Having $f,f^{\prime}\in F$ apply the algorithm of enumeration of the set $\mathsf{Wrd}^{=}_{\nu}$ for verification if $\nu(n_{i}*f)=\nu(f^{\prime})$, $i=1,2$. When we get a confirmation of this equality, we extend the corresponding $\Sigma_{i}$ by $(f,f^{\prime})$. We apply it simultaneously to each pair $(f,f^{\prime})$. Since
$\forall n\in E\,\,(|\nu(F)\cap\nu(n)\nu(F)|\geq\frac{2}{3}|\nu(F)|)$
@@ 4.1 Algorithmic complexity of computable Følner sequences
Given number $s$ the inequality $\frac{|F\setminus x*F|}{|F|}<\frac{1}{n}$ can be verified effectively. Since the set of numbers of all total functions belongs to the class $\Sigma_{2}^{0}$ (see [34]), it is easy to see that the set of all $m$ which satisfy (9) is a $\Pi_{3}^{0}$ set. This proves the first part of the theorem.
@@
For a given $s$, we use the enumeration of the set $\{(n,x):x\in W_{\varphi_{e}(n)}\}$ to find the element $(n_{s},x_{s})$ with the number $s$. For each $i=1,\ldots,s$ such that $i\neq n_{s}$ let $F_{s,i}=\{0,g^{\pm 1}_{i},g_{i}^{\pm 2},\ldots,g_{i}^{\pm s}\}$. For $i=n_{s}$ we put $F_{s,i}=\{0,g^{\pm 1}_{i}\}$. Let $F_{s}^{e}=\bigoplus\limits_{1}^{s}F_{s,i}$. Then in the former case $F_{s}^{e}$ is an $\frac{1}{2s}$-Følner set with respect to $g^{\pm 1}_{i}$ and in the latter case $F_{s}^{e}$ is not a $\frac{1}{4}$-Følner set with respect to $g_{i}$. This ends the construction. Note that it depends on $e$ and corresponds to a Turing machime which can be explicitly found.
@@
Since for every $e$ the number of the algorithm producing $\{F^{e}_{s}\}$ can be effectively found it follows that the set $\overline{Cof}$ is reducible to $\mathfrak{F}_{seq}(G)$, which completes the proof. ∎
@@ 4.2 Complexity of convergence moduli
It is easy to see that the complement of $Mod_{\mathcal{F}}({\bf x})$ in $\mathbb{N}\times\mathbb{N}$ is computably enumerable. Furthermore, when $Mod_{\mathcal{F}}({\bf x})$ is computable, the function
@@
(As we already noted in Section 2, $\{-i,-i+1,\ldots,0,\ldots,i-1,i\}$ is $\frac{1}{2i}$-Følner with respect to the generator of $\mathbb{Z}$.) Our main result below shows that the growth of $mod_{\mathcal{F},{\bf x}}$ is not bounded by a primitive recursive function. We believe that our main construction can be adapted to many other computable groups. The construction is presented in the proof of the following theorem.
@@
Let $f$ be a total computable function $\mathbb{N}\to\mathbb{N}$. Then there is a computable ${\bf x}_{0}\in 2^{\mathbb{Z}}$ such that with respect to the computable Følner sequence $\mathcal{F}$ we have $m_{i}({\bf x}_{0})\to 0$, $i\in\mathbb{N}$, and for every $k\in\mathbb{N}$ there is $j>f(k)$ such that $|m_{j}({\bf x}_{0})|\geq\frac{1}{k}$.
@@
We start at $k=3$. At inductive step $k$ assume that at step k1>1k-1>1 we have already defined some number $i_{k-1}$ and values ${\bf x}_{0}(i)$ for all $i$ with $i_{k-1}\leq i\leq i_{k-1}$. We also assume that
@@
To have this condition at step k1=2k-1=2 we may assume that $i_{2}=5$ , ${\bf x}_{0}(1)={\bf x}_{0}(2)={\bf x}_{0}(3)=0$ and ${\bf x}_{0}(4)={\bf x}_{0}(5)=1$. Note that it now implies
@@
There is a computable ${\bf x}_{0}\in 2^{\mathbb{Z}}$ such that with respect to the computable Følner sequence $\mathcal{F}$ we have $m_{i}({\bf x}_{0})\to 0$, $i\in\mathbb{N}$, and for every primitive recursive function $f$ there is $k\in\mathbb{N}$ and $j>f(k)$ such that $|m_{j}({\bf x}_{0})|\geq\frac{1}{k}$.
@@
Since this function has finitely many values, according to Lemma 2.1 from [23] and example (iii) after it, it is a partial recursive function from $[0,1]$ to $[0,1]$. At this stage we remind the reader that according to Definition 2.1 of [23] intuitively the computation of a partial recursive real function $f$ is as follows. For given dyadic number $\bf x$} and and a natural number $n$, the Turing machine tries to find a dyadic rational number $d_{n}$ of length $n$ such that $d_{n}$ is close to $f({\bf x})$ up to $\frac{1}{2^{n}}$. During the computation, $\bf x$} is used as an oracle. As a result $f({\bf x})=\lim_{n}d_{n}$.
@@
$\bullet$ for every interval $(q,q^{\prime})\subseteq[0,1]$, for every real $\varepsilon>0$ and for every set $X\subseteq G$ that is represented by some $r\in(q,q^{\prime})$ (as an element of $2^{G}$) there is $X^{\prime}\subseteq G$ and $r^{\prime}\in(q,q^{\prime})$ which represents $X^{\prime}$ such that $|r-r^{\prime}|<\varepsilon$ but $|m(r)-m(r^{\prime})|>\frac{1}{2}$.

### v2

@@ Abstract
The paper considers computable Følner sequences in computably enumerable amenable groups. We extend some basic results of M. Cavaleri on existence of such sequences to the case of groups where finite generation is not assumed. We also initiate some new directions in this topic, for example complexity of families of effective Følner sequences. Possible extensions of this approach to metric groups are also discussed.
@@ 1 Introduction
Since being finitely generated is not necessary for amenability, the question arises what happens if we consider the case of recursively presented groups without the assumption of finite generation. According to the approach of computable algebra, the question concerns the class of computably enumerable numbered groups and the subclass of computable numbered groups, a counterpart of decidability of the word problem. These notions are thoroughly discribed in Section 2. The following theorem generalizes aforementioned results of Cavaleri to the case of computably enumerable numbered groups.
@@
Let $(G,\nu)$ be a computably enumerable numbered group. The following conditions are equivalent:
@@
We also compare convergence moduli of sequences of means corresponding to these Følner sequences. In particular we show that in the case of the standard Følner sequence of $(\mathbb{Z},+)$ and the corresponding sequence of means $m_{i}({\bf x})$ (which converge to an invariant mean witnessing amenability of $\mathbb{Z}$) the following statement holds.
Theorem 3.
For any total computable $f:\mathbb{N}\to\mathbb{N}$ there is a computable ${\bf x}_{0}\in 2^{\mathbb{Z}}$ such that the seqience $m_{i}({\bf x}_{0})$, $i\in\mathbb{N}$, converges to 0, but for every $k\in\mathbb{N}$ there is $j>f(k)$ such that $|m_{j}({\bf x}_{0})|\geq\frac{1}{k}$.
@@
The authors are grateful to M. Cavaleri and T. Ceccherini-Silberstein for reading of some preliminary versions of the paper and helpful remarks.
@@ 2 Preliminaries
Concerning other details, let us mention that we often identify finite sets $F\subset\mathbb{N}$ with their Gödel numbers. For any sets $X$ and $Y$ we will write $X\subset_{fin}Y$ to denote that $X$ is a finite subset of $Y$. Throughout this paper a group (metric group) $G$ is a countable/separable group without any presumption about its generating set.
@@ 2.1 Computabie presentations
We emphasize that this group is computable, i.e. its operations are computable functions! For every recursively presented group $G=\langle X\rangle$ and a natural homomorphism $\rho:\mathbb{F}_{\omega}\rightarrow G$ (taking $\omega$ onto $X$) we obtain that $\rho$ is a numbering of a computably enumerable presentation of $G$. This argument can be generalized as follows.
@@
    The set MultT := {(i,j,k):  ν(i)ν(j)=ν(k)} is computable if and only if the set $\{(n_{1},n_{2}):\nu(n_{1})=\nu(n_{2})\}$ is computable;
@@
From now on we consider computably enumerable groups in the following way.
  • Any computably enumerable group $(G,\nu)$ is taken with a homomorphis $\nu:(\mathbb{N},\star,^{-1},1)\to G$ where $(\mathbb{N},\star,^{-1},1)$ is a group with computable operations, i.e. satisfying the conditions of Lemma 2.3.
  • If $(G,\nu)$ is computable, we additionally assume that $\nu$ is an isomorphism.
@@ 2.2 Amenability
$\forall x\in D\hskip 14.22636pt\frac{|F\setminus xF|}{|F|}\leq\frac{1}{n}.$ (1)
@@
A sequence $(F_{j})_{j\in\mathbb{N}}$ of non-empty finite subsets of $G$ is a (left) Følner sequence, if for every $g\in G$ the following condition holds:
@@
Suppose that $G$ satisfies the Følner conditions. Let $(F_{j}\,|\,j\in J)$ be a left Følner sequence of $G$. Consider, for each $j\in J$, the mean with finite support $m_{j}:\ell^{\infty}(G)\to\mathbb{R}$ defined by
@@
The following facts are well known and/or easy to prove (see also Remark 2.2 from [3]).
@@ 3 Effective amenability of computably enumerable groups
In this section $(G,\nu)$ is a numbered metric group such that $\nu$ is a homomorphism from $(\mathbb{N},\star,^{-1},1)$ to $G$ where operations $\star$ and -1 are computable. In the case of Følner’s condition of amenability, we consider two types of effectiveness.
@@
Some variants of these notions were characterized by M. Cavaleri in [3] in the case of finitely generated groups. The goal of this section is an adaptation of these characterizations in our general case. Although we use the same arguments, the adaption needs some additional effort.
@@ 3.1 Computable Reiter’s functions
The main result of this part, Theorem 3.5, is a natural generalization of a theorem of M. Cavaleri from [3] (Theorem 3.1) to the case of groups which are not finitely generated. Throughout this section we assume that $(G,\nu)$ is a computably enumerable group under a homomrphism $\nu$ as in the beginning of the section.
The following notation will be used below. It corresponds to Section 3 of [3]. If $f:\mathbb{N}\to\mathbb{R}^{+}$ is summable and $g\in G$, then let
@@
Having $x\in\mathbb{N}$ and an equivalence relation $E$ on the set $F\cup x^{-1}\star F$ fix a set $F_{0}\subset F$ of representatives of the classes of $E$ in $F$. Then we define the positive rational number:
$M_{E,F}^{x}(f):=\frac{\sum\{|(E_{F,1}(f)(v)-E_{F,1}(f)(x^{-1}\star v))|\,|\,v\in F_{0}\}}{\sum_{v\in F}f(v)}.$
We denote by $P$ the canonical equivalence relation of the set $F\cup x^{-1}\star F$, i.e. the partition into sets $\{\nu^{-1}(\nu(k))\,|\,k\in F\cup x^{-1}\star F\}$. Then for every $x\in\mathbb{N}$ we have
@@
For any two equivalence relations $E$ and $E^{\prime}$ of the set $F\cup x^{-1}\star F$ with $E\subseteq E^{\prime}$, the triangle inequality implies $M_{E,F}^{x}(f)\geq M_{E^{\prime},F}^{x}(f)$. In particular, for any $x\in\mathbb{N}$ and $E\subseteq P$ on $F\cup x^{-1}\star F$ we have:
@@
The algorithm $\varkappa(n,D,f)$ starts as follows. For an input $f$ let $F=\mathsf{supp}(f)$. Put $P_{0}$ to be the (finest) partition of $\bigcup_{g\in D}g^{-1}\star F$ into singletons. At the $m$-th step of the enumeration of $\{(n_{1},n_{2}):\nu(n_{1})=\nu(n_{2})\}$ we are trying to merge classes of the equivalence relation $P_{m-1}$ on $\bigcup_{g\in D}g^{-1}\star F$ already constructed at step m1m-1. We do so when we meet $(n_{1},n_{2})\not\in P_{m-1}$ with $\nu(n_{1})=\nu(n_{2})$. In this case we just merge the classes of $n_{1}$ and $n_{2}$. In this way we obtain $P_{m}\subseteq P$. Then we verify if $M_{P_{m},F}^{x}(f)\leq\frac{1}{n}$ for all $x\in D$. We stop $\varkappa(n,D,f)$ when these inequalities hold. In this case by (5) and (6), the function $\nu_{G,1}(f)$ is $\frac{1}{n}$-invariant with respect to $D$. If there exist $x$, such that $M_{P_{m},F}^{x}(f)>\frac{1}{n}$ and $P_{m}=P$, then the function $\nu_{G,1}(f)$ is not $\frac{1}{n}$-invariant. Note that there is no algorithm for recognizing the latter possibility.
@@ 3.2 Effective amenability of computable groups
Fix $n_{1},n_{2}$. Let $E$ be the set $\{n_{1},n_{2}\}$. We use the algorithm for (ii) to find a set $F$ corresponding to $5$ and $E$, i.e. $\nu(F)\in F\o l_{G,\nu(E)}(5)$ and $|F|=|\nu(F)|$. Let $F=\{f_{1},f_{2},\ldots,f_{k}\}$.
For each $i\in\{1,2\}$ we define $\Sigma_{i}\subseteq\{(f,f^{\prime})\,|\,\nu(n_{i})\nu(f)=\nu(f^{\prime})\,,\,f,f^{\prime}\in F\}$ by the following procedure. Having $f,f^{\prime}\in F$ apply the algorithm of enumeration of the set $\mathsf{Wrd}^{=}_{\nu}$ for verification if $\nu(n_{i}\star f)=\nu(f^{\prime})$, $i=1,2$. When we get a confirmation of this equality, we extend the corresponding $\Sigma_{i}$ by $(f,f^{\prime})$. We apply it simultaneously to each pair $(f,f^{\prime})$. Since
$\forall n\in E\,\,(|\nu(F)\cap\nu(n)\nu(F)|\geq\frac{3}{5}|\nu(F)|)$
@@ 4.1 Algorithmic complexity of computable Følner sequences
Given number $s$ the inequality $\frac{|F\setminus x\star F|}{|F|}<\frac{1}{n}$ can be verified effectively. Since the set of numbers of all total functions belongs to the class $\Sigma_{2}^{0}$ (see [33]), it is easy to see that the set of all $m$ which satisfy (7) is a $\Pi_{3}^{0}$ set. This proves the first part of the theorem.
@@
For a given $s$, we use the enumeration of the set $\{(n,x):x\in W_{\varphi_{e}(n)}\}$ to find the element $(n_{s},x_{s})$ with the number $s$. For each $i=1,\ldots,s$ such that $i\neq n_{s}$ let $F_{s,i}=\{0,g^{\pm 1}_{i},g_{i}^{\pm 2},\ldots,g_{i}^{\pm s}\}$. For $i=n_{s}$ we put $F_{s,i}=\{0,g^{\pm 1}_{i}\}$. Then in the former case $F_{s}^{e}$ is an $\frac{1}{2s}$-Følner set with respect to $g^{\pm 1}_{i}$, and in the latter case $F_{s}^{e}$ is not a $\frac{1}{4}$-Følner set with respect to $g_{i}$. Let $F_{s}^{e}=\bigoplus\limits_{1}^{s}F_{s,i}$. This ends the construction. Note that it depends on $e$, and is realized by a Turing machime; the latter can be explicitly found.
@@
Since for every $e$ the number of the algorithm producing $\{F^{e}_{s}\}$ can be effectively found, it follows that the set $\overline{Cof}$ is reducible to $\mathfrak{F}_{seq}(G)$, which completes the proof. ∎
@@ 4.2 Complexity of convergence moduli
It is easy to see that the complement of $Mod_{\mathcal{F}}({\bf x})$ in $\mathbb{N}\times\mathbb{N}$ is computably enumerable. Furthermore, when $Mod_{\mathcal{F}}({\bf x})$ is computably enumerable, the function
@@
(As we already noted in Section 2, $\{-i,-i+1,\ldots,0,\ldots,i-1,i\}$ is $\frac{1}{2i}$-Følner with respect to the generator of $\mathbb{Z}$.) Our main result below shows that the growth of $mod_{\mathcal{F},{\bf x}}$ is not bounded by a primitive recursive function. We believe that our main construction can be adapted to many other computable groups. The construction is presented in the proof of the following theorem (Theorem 3 from Introduction).
@@
Let $f$ be a total computable function $\mathbb{N}\to\mathbb{N}$. Then there is a computable ${\bf x}_{0}\in 2^{\mathbb{Z}}$ such that with respect to the computable Følner sequence $\mathcal{F}$ the values $m_{i}({\bf x}_{0})$, $i\in\mathbb{N}$, converge to 0, and for every $k\in\mathbb{N}$ there is $j>f(k)$ such that $|m_{j}({\bf x}_{0})|\geq\frac{1}{k}$.
@@
At inductive step $k$ assume that at step k1>1k-1>1 we have already defined some number $i_{k-1}$ and values ${\bf x}_{0}(i)$ for all $i$ with $i_{k-1}\leq i\leq i_{k-1}$. We start at $k=3$, where it is assumed that $i_{2}=5$ , ${\bf x}_{0}(1)={\bf x}_{0}(2)={\bf x}_{0}(3)=0$ and ${\bf x}_{0}(4)={\bf x}_{0}(5)=1$. At every step $k>2$ we also assume that
@@
It is easy to see from our description of step 3 that this condition is satisfied for k1=2k-1=2. Note that it implies
@@
There is a computable ${\bf x}_{0}\in 2^{\mathbb{Z}}$ such that with respect to the computable Følner sequence $\mathcal{F}$ the values $m_{i}({\bf x}_{0})$, $i\in\mathbb{N}$, converge to 0, and for every primitive recursive function $f$ there is $k\in\mathbb{N}$ and $j>f(k)$ such that $|m_{j}({\bf x}_{0})|\geq\frac{1}{k}$.
@@
Since this function has finitely many values, according to Lemma 2.1 from [22] and example (iii) after it, it is a partial recursive function from $[0,1]$ to $[0,1]$. At this stage we remind the reader that according to Definition 2.1 of [22] intuitively the computation of a partial recursive real function $f$ is as follows. For a given dyadic number $\bf x$} and and a natural number $n$, the Turing machine tries to find a dyadic rational number $d_{n}$ of length $n$ such that $d_{n}$ is close to $f({\bf x})$ up to $\frac{1}{2^{n}}$. During the computation, $\bf x$} is used as an oracle. As a result $f({\bf x})=\lim_{n}d_{n}$.
@@
$\bullet$ for every interval $(q,q^{\prime})\subseteq[0,1]$, for every real $\varepsilon>0$ and for every set $X\subseteq G$ that is represented by some $r\in(q,q^{\prime})$ (as an element of $2^{G}$), there is $X^{\prime}\subseteq G$ and $r^{\prime}\in(q,q^{\prime})$ which represents $X^{\prime}$ such that $|r-r^{\prime}|<\varepsilon$ but $|m(r)-m(r^{\prime})|>\frac{1}{2}$.